Determine the total angular momentum quantum number if the minimum angle between the total angular momentum vector (in a single electron atom) and the z axis is 28.13°.
1
Expert's answer
2020-05-28T12:01:29-0400

The minimum angle between the  total angular momentum vector (in a single electron atom) and the z axis is 

"cos\theta=\frac{l (\frac{h}{2\pi})}{\sqrt{l(l+1)}(\frac{h}{2\pi})}"

"cos28.13°=\frac{l (\frac{h}{2\pi})}{\sqrt{l(l+1)}(\frac{h}{2\pi})}"

"cos28.13°=\frac{l }{\sqrt{l(l+1)}}"

"0.88=\frac{l }{\sqrt{l(l+1)}}"

on solving the above equation, we get "l=3.43=3 (Approximately)"

So the total angular momentum is

"L=\sqrt{l(l+1)}(\frac{h}{2\pi})=\sqrt{3(3+1)}(\frac{h}{2\pi})=\sqrt{12}(\frac{h}{2\pi})"

"L=\sqrt{12}(\frac{h}{2\pi})=\sqrt{12}(\frac{6.626*10^{-34}}{2\pi})=3.65*10^{-34} J.s"